# Nearly-Linear Time Seeded Extractors with Short Seeds

### v1

@@ Abstract
Seeded extractors are fundamental objects in pseudorandomness and cryptography, and a deep line of work has designed polynomial-time seeded extractors with nearly-optimal parameters. However, existing constructions of seeded extractors with short seed length and large output length run in time $\Omega(n\log(1/\varepsilon))$ and often slower, where $n$ is the input source length and $\varepsilon$ is the error of the extractor. Since cryptographic applications of extractors require $\varepsilon$ to be small, the resulting runtime makes these extractors unusable in practice.
Motivated by this, we explore constructions of strong seeded extractors with short seeds computable in nearly-linear time $O(n\log^{c}n)$, for any error $\varepsilon$. We show that an appropriate combination of modern condensers and classical approaches for constructing seeded extractors for high min-entropy sources yields strong extractors for $n$-bit sources with any min-entropy $k$ and any target error $\varepsilon$ with seed length $d=O(\log(n/\varepsilon))$ and output length $m=(1-\eta)k$ for an arbitrarily small constant $\eta>0$, running in nearly-linear time, after a reasonable one-time preprocessing step (finding a primitive element of $\mathds{F}_{q}$ with $q=\operatorname{poly}(n/\varepsilon)$ a power of $2$) that is only required when $k<2^{C\log^{*}\!n}\cdot\log^{2}(n/\varepsilon)$, for a constant $C>0$ and $\log^{*}\!$ the iterated logarithm, and which can be implemented in time $\operatorname{polylog}(n/\varepsilon)$ under mild conditions on $q$. As a second contribution, we give an instantiation of Trevisan’s extractor that can be evaluated in truly linear time in the RAM model, as long as the number of output bits is at most $\frac{n}{\log(1/\varepsilon)\operatorname{polylog}(n)}$. Previous fast implementations of Trevisan’s extractor ran in $\widetilde{O}(n)$ time in this setting. In particular, these extractors directly yield privacy amplification protocols with the same time complexity and output length, and communication complexity equal to their seed length.
@@ 1 Introduction
Seeded randomness extractors are central objects in the theory of pseudorandomness. A strong $(k,\varepsilon)$-seeded extractor is a deterministic function $\mathsf{Ext}\colon\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{n}\times\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{d}\to\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{m}$ that receives as input an $n$-bit source of randomness $X$ with $k$ bits of min-entropy [^1] A random variable $X$ has $k$ bits of min-entropy if $\Pr[X=x]\leq 2^{-k}$ for all $x$. Min-entropy has been the most common measure for the quality of a weak source of randomness since the work of Chor and Goldreich [CG88]. and a $d$-bit independent and uniformly random seed $Y$, and outputs an $m$-bit string $\mathsf{Ext}(X,Y)$ that is $\varepsilon$-close in statistical distance to the uniform distribution over $\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{m}$, where $\varepsilon$ is an error term, even when the seed $Y$ is revealed. Besides their most direct application to the generation of nearly-perfect randomness from imperfect physical sources of randomness (and their inaugural applications to derandomizing space-bounded computation [NZ96] and privacy amplification [BBCM95]), seeded extractors have also found many other surprising applications throughout computer science, particularly in cryptography.
@@
For most applications, it is important to minimize the seed length of the extractor. A standard application of the probabilistic method shows the existence of strong $(k,\varepsilon)$-seeded extractors with seed length $d=\log(n-k)+2\log(1/\varepsilon)+O(1)$ and output length $m=k-2\log(1/\varepsilon)-O(1)$, and we also know that these parameters are optimal up to the $O(1)$ terms [RT00]. This motivated a deep line of research devising explicit constructions of seeded extractors with seed length as small as possible spanning more than a decade (e.g., [NZ96, SZ99, NT99, Tre01, TZS06, SU05]) and culminating in extractors with essentially optimal seed length [LRVW03, GUV09]. In particular, the beautiful work of Guruswami, Umans, and Vadhan [GUV09] gives explicit strong extractors with order-optimal seed length $d=O(\log(n/\varepsilon))$ and output length $m=(1-\eta)k$ for any constant $\delta>0$, and follow-up work [DKSS13, TU12] further improved the entropy loss $k+d-m$. The extractors constructed in these works are explicit, in the sense that there is an algorithm that given $x$ and $y$ computes the corresponding output $\mathsf{Ext}(x,y)$ in time polynomial in the input length.
@@
A closer look shows that the short-seed constructions presented in the literature all run in time $\Omega(n\log(1/\varepsilon))$, and often significantly slower. In cryptographic applications of extractors we want the error guarantee $\varepsilon$ to be small, which means that implementations running in time $\Omega(n\log(1/\varepsilon))$ are often impractical. If we insist on nearly-linear runtime for arbitrary error $\varepsilon$, we can use strong seeded extractors based on universal hash functions that can be implemented in $O(n\log n)$ time (e.g., see [HT16]), have essentially optimal output length, but have the severe drawback of requiring a very large seed length $d=\Omega(m)$.
@@
These limitations have been noted in a series of works studying concrete implementations of seeded extractors, with practical applications in quantum cryptography in mind [MPS12, FWE+23, FYEC24]. For example, Foreman, Yeung, Edgington, and Curchod [FYEC24] implement a version of Trevisan’s extractor [Tre01, RRV02] with its standard instantiation of Reed–Solomon codes concatenated with the Hadmadard code, and emphasize its excessive running time as a major reason towards non-adoption. [^2] The reason why these works focus on Trevisan’s extractor is that this is the best seeded extractor (in terms of asymptotic seed length) that is known to be secure against quantum adversaries [DPVR12]. Instead, they have to rely on extractors based on universal hash functions, which, as mentioned above, are fast but require very large seeds.
@@
Progress on this problem would immediately lead to faster implementations of many cryptographic protocols that use seeded extractors.
@@ 1.1 Our Contributions
We make progress on the construction of nearly-linear time extractors.
@@
We construct nearly-linear time strong seeded extractors with order-optimal seed length and large output length for any $k$ and $\varepsilon$, with the caveat that they require a one-time preprocessing step whenever $k=O(\log^{2}(n/\varepsilon))$. This preprocessing step corresponds to finding primitive elements of finite fields $\mathds{F}_{q}$ with $q=\operatorname{poly}(n/\varepsilon)$, which, as we discuss below, is reasonable in practical applications. More precisely, we have the following result.
@@
  • if $k\geq 2^{C\log^{*}\!n}\cdot\log^{2}(n/\varepsilon)$, then $\mathsf{Ext}$ is computable in time $\widetilde{O}(n)$, where $\widetilde{O}(\cdot)$ hides polylogarithmic factors in its argument and $\log^{*}\!$ denotes the iterated logarithm;
  • if $k<2^{C\log^{*}\!n}\cdot\log^{2}(n/\varepsilon)$, then $\mathsf{Ext}$ is computable in time $\widetilde{O}(n)$ after a preprocessing step, corresponding to finding a primitive element of $\mathds{F}_{q}$ with $q=\operatorname{poly}(n/\varepsilon)$ a power of $2$. [^3] In full rigor, the preprocessing step corresponds to finding primitive elements of $O(\log\log n)$ fields $\mathds{F}_{q}$ with orders $q\leq\operatorname{poly}(n/\varepsilon)$, each a power of $2$. This $O(\log\log n)$ term has negligible influence on the complexity of this preprocessing step. Note that we can find such a primitive element in time $\operatorname{polylog}(n/\varepsilon)$ if $q\leq\operatorname{poly}(n/\varepsilon)$ is a power of $2$ and we know the factorization of $q-1$, but we do not know how to do that in time $\widetilde{O}(\log q)$. More precisely, given the factorization of $q-1$ we can test whether a given $\alpha\in\mathds{F}_{q}$ is primitive in time $\operatorname{polylog}(q)$ by checking whether $\alpha^{\frac{q-1}{p}}\neq 1$ for all prime factors $p$ of $q-1$. We can exploit this in various ways. If we are fine with using randomness in the one-time preprocessing stage, then we can sample an element of $\mathds{F}_{q}$ uniformly at random, test whether it is primitive, and repeat if not. If we insist on a deterministic algorithm, then we can combine the testing procedure with algorithms of Shoup [Sho90] or Shparlinski [Shp92] which identify in time $\operatorname{polylog}(q)$ a subset of size $\operatorname{polylog}(q)$ in $\mathds{F}_{q}$ that is guaranteed to contain a primitive element. For an alternative faster randomized algorithm, see [DD06].
1 follows from combining modern condensers with short seeds (namely, the lossless condenser of Kalev and Ta-Shma [KT22] and the lossy Reed-Solomon-based condenser of Guruswami, Umans, and Vadhan [GUV09]) with a careful combination and instantiation of classical recursive approaches developed by Srinivasan and Zuckerman [SZ99] and in [GUV09]. It readily implies, among other things, an $\widetilde{O}(n)$-time privacy amplification protocol where only $O(\log(n/\varepsilon))$ bits need to be communicated over the one-way authenticated public channel and almost all the min-entropy can be extracted (after a reasonable one-time preprocessing step if the min-entropy bound $k$ is very small).
@@
As a conceptual contribution which may be of independent interest, we present a new “non-recursive” construction of extractors with seed length $O(\log(n/\varepsilon))$ and output length $(1-\eta)k$ that is computable in nearly-linear time when $k>\operatorname{polylog}(1/\varepsilon)$ and avoids the complicated recursive procedures from [SZ99, GUV09]. We believe this to be a conceptually better approach towards constructing seeded extractors, and we discuss it in more detail in the technical overview.
@@
One of the most widely-used explicit seeded extractors is Trevisan’s extractor [Tre01, RRV02]. While by now we have extractors with better parameters, one of its main advantages is that it is one of the few examples of extractors, and in a sense the best one, which are known to be quantum proof. [^4] An extractor is quantum proof if its output is close to uniform even in the presence of a quantum adversary that has some (bounded) correlation with $X$. A bit more formally, $\mathsf{Ext}$ is quantum-proof if for all classical-quantum state $\rho_{XE}$ (where $E$ is a quantum state correlated with $X$) with $H_{\infty}(X|E)\geq k$, and a uniform seed $Y$, it holds that $\rho_{\mathsf{Ext}(X,Y)YE}\approx_{\varepsilon}\rho_{U_{m}}\otimes\rho_{Y}\otimes\rho_{E}$. See [DPVR12] for more details.
Trevisan’s extractor uses two basic primitives: combinatorial designs (when more than one output bit is desired), and binary list-decodable codes. A standard instantiation of such suitable codes goes by concatenating a Reed-Solomon code with a Hadamard code, and this is also what is considered in [FWE+23, FYEC24]. As they also observe, this gives a nearly-linear time construction when the output length $m=1$. In fact, by leveraging fast multipoint evaluation, one can also get a nearly-linear time construction for any output length $m\leq\frac{n}{\log(1/\varepsilon)}$, although this was not noted in previous works. [^5] For a rigorous statement on fast multipoint evaluation, see Lemma 2.1.
Our main contribution in this direction is an alternative instantiation of Trevisan’s extractor that can be computed in truly linear time on a RAM in the logarithmic cost model, for any output length $m\leq\frac{n}{\log(1/\varepsilon)\cdot\operatorname{polylog}(n)}$.
@@
    Time $O(n)+m\log(1/\varepsilon)\cdot\operatorname{polylog}(n)$ after a preprocessing step running in time $\widetilde{O}(m\log(n/\varepsilon))$, on a RAM in the logarithmic cost model. In particular, there exists a universal constant $c$, such that whenever $m\leq\frac{n}{\log(1/\varepsilon)\cdot\log^{c}(n)}$, the instantiation runs in time $O(n)$, without the need for a preprocessing step.
@@ 1.2 Other Related Work
Besides the long line of work focusing on improved constructions of explicit seeded extractors and mentioned in the introduction above, other works have studied randomness extraction in a variety of restricted computational models. These include extractors computable by streaming algorithms [BRST02], local algorithms [Lu02, Vad04, BG13, CL18], AC0 circuits [GVW15, CL18, CW24], AC0 circuits with a layer of parity gates [HIV22], NC1 circuits [CW24], and low-degree polynomials [ACG+22, AGMR24, GGH+24]. Moreover, implementations in various restricted computational models of other fundamental pseudorandomness primitives such as $k$-wise and $\varepsilon$-biased generators, that often play a key role in constructions of various types of extractors, have also been independently studied (see [HV06, Hea08, CRSW13, MRRR14] for a very partial list).
As mentioned briefly above, some works have also focused on constructing seeded extractors computable in time $O(n\log n)$ motivated by applications in privacy amplification for quantum key distribution. Such constructions are based on hash functions, and are thus far restricted to $\Omega(m)$ seed length. The work of Hayashi and Tsurumaru [HT16] presents an extensive discussion of such efforts. We also mention that nearly-linear time extractors with very short seed, in the regime $k=n^{\Omega(1)}$ and $\varepsilon=n^{-o(1)}$, were given in [DMOZ22], with applications in derandomization.

### v2

@@ Abstract
Seeded extractors are fundamental objects in pseudorandomness and cryptography, and a deep line of work has designed polynomial-time seeded extractors with nearly-optimal parameters. However, existing constructions of seeded extractors with short seed length and large output length run in time $\Omega(n\log(1/\varepsilon))$ and often slower, where $n$ is the input source length and $\varepsilon$ is the error of the extractor. Since cryptographic applications of extractors require $\varepsilon$ to be small, the resulting runtime makes these extractors impractical.
Motivated by this, we explore constructions of strong seeded extractors with short seeds computable in nearly-linear time $O(n\log^{c}n)$, for any error $\varepsilon$. We show that an appropriate combination of modern condensers and classical approaches for constructing seeded extractors for high min-entropy sources yields such extractors. More precisely, we obtain strong extractors for $n$-bit sources with any min-entropy $k$ and any target error $\varepsilon$ with seed length $d=O(\log(n/\varepsilon))$ and output length $m=(1-\eta)k$ for an arbitrarily small constant $\eta>0$, running in nearly-linear time. When $k$ or $\varepsilon$ are very small, our construction requires a reasonable one-time preprocessing step. These extractors directly yield privacy amplification protocols with nearly-linear time complexity (possibly after a one-time preprocessing step), large output length, and low communication complexity. As a second contribution, we give an instantiation of Trevisan’s extractor that can be evaluated in truly linear time in the RAM model, as long as the number of output bits is at most $\frac{n}{\log(1/\varepsilon)\operatorname{polylog}(n)}$. Previous fast implementations of Trevisan’s extractor ran in $\widetilde{O}(n)$ time in this setting.
@@ 1 Introduction
Seeded randomness extractors are central objects in the theory of pseudorandomness. A strong $(k,\varepsilon)$-seeded extractor, first introduced by Nisan and Zuckerman [NZ96], is a deterministic function $\mathsf{Ext}\colon\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{n}\times\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{d}\to\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{m}$ that receives as input an $n$-bit source of randomness $X$ with $k$ bits of min-entropy [^1] A random variable $X$ has $k$ bits of min-entropy if $\Pr[X=x]\leq 2^{-k}$ for all $x$. Min-entropy has been the most common measure for the quality of a weak source of randomness since the work of Chor and Goldreich [CG88]. and a $d$-bit independent and uniformly random seed $Y$, and outputs an $m$-bit string $\mathsf{Ext}(X,Y)$ that is $\varepsilon$-close in statistical distance to the uniform distribution over $\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{m}$, where $\varepsilon$ is an error term, even when the seed $Y$ is revealed. Besides their most direct application to the generation of nearly-perfect randomness from imperfect physical sources of randomness (and their early applications to derandomizing space-bounded computation [NZ96] and privacy amplification [BBR88, BBC+95]), seeded extractors have also found many other surprising applications throughout computer science, particularly in cryptography (specifically, in leakage-resilient cryptography [SV19, QWW21] and non-malleable cryptography [BGW19, CGL20, AKO+22]).
@@
For most applications, it is important to minimize the seed length of the extractor. A standard application of the probabilistic method shows the existence of strong $(k,\varepsilon)$-seeded extractors with seed length $d=\log(n-k)+2\log(1/\varepsilon)+O(1)$ and output length $m=k-2\log(1/\varepsilon)-O(1)$, and we also know that these parameters are optimal up to the $O(1)$ terms [RT00]. This motivated a deep line of research devising explicit constructions of seeded extractors with seed length as small as possible spanning more than a decade (e.g., [NZ96, SZ99, NT99, TRE01, TZS06, SU05]) and culminating in extractors with essentially optimal seed length [LRV+03, GUV09]. In particular, the beautiful work of Guruswami, Umans, and Vadhan [GUV09] gives explicit strong extractors with order-optimal seed length $d=O(\log(n/\varepsilon))$ and output length $m=(1-\eta)k$ for any constant $\eta>0$, and follow-up work [DKS+13, TU12] further improved $m$ to (1o(1))k(1-o(1))k (at the expense of higher error). The extractors constructed in these works are explicit, in the sense that there is an algorithm that given $x$ and $y$ computes the corresponding output $\mathsf{Ext}(x,y)$ in time polynomial in the input length.
@@
A closer look shows that the short-seed constructions presented in the literature all run in time $\Omega(n\log(1/\varepsilon))$, and often significantly slower. In cryptographic applications of extractors we want the error guarantee $\varepsilon$ to be small, which means that implementations running in time $\Omega(n\log(1/\varepsilon))$ are often impractical. If we insist on nearly-linear runtime for arbitrary error $\varepsilon$, we can use strong seeded extractors based on universal hash functions that can be implemented in $O(n\log n)$ time (e.g., see [HT16]) and have essentially optimal output length, but have the severe drawback of requiring a very large seed length $d=\Omega(m)$.
@@
These limitations have been noted in a series of works studying concrete implementations of seeded extractors, with practical applications in quantum cryptography in mind [MPS12, FWE+23, FYE+25]. For example, Foreman, Yeung, Edgington, and Curchod [FYE+25] implement a version of Trevisan’s extractor [TRE01, RRV02] with its standard instantiation of Reed–Solomon codes concatenated with the Hadmadard code, and emphasize its excessive running time as a major reason towards non-adoption. [^2] The reason why these works focus on Trevisan’s extractor is that this is the best seeded extractor (in terms of asymptotic seed length) that is known to be secure against quantum adversaries [DPV+12]. Instead, they have to rely on extractors based on universal hash functions, which, as mentioned above, are fast but require very large seeds.
@@
Progress on this problem would immediately lead to faster implementations of many cryptographic protocols that use seeded extractors, like those mentioned above – the most significant performance gains would be in the context of privacy amplification.
@@ 1.1 Our Contributions
We make progress on the construction of nearly-linear time seeded extractors.
@@
We construct nearly-linear time strong seeded extractors with order-optimal seed length and large output length for any $k$ and $\varepsilon$, with the caveat that they require a reasonable one-time preprocessing step whenever $k$ or $\varepsilon$ are small. More precisely, we have the following result.
@@
  1. 1.
[…]
  2. 2.
    If $k\geq 2^{C\log^{*}\!n}\cdot\log^{2}(n/\varepsilon)$ and $\varepsilon<2^{-Cn^{0.1}}$, then $\mathsf{Ext}$ is computable in time $\widetilde{O}(n)$ after a preprocessing step, corresponding to generating $O(\log^{*}\!n)$ primes $q\leq\operatorname{poly}(n/\varepsilon)$;
  3. 3.
    If $k<2^{C\log^{*}\!n}\cdot\log^{2}(n/\varepsilon)$, then $\mathsf{Ext}$ is computable in time $\widetilde{O}(n)$ after a preprocessing step, corresponding to generating $O(\log\log n)$ primes $q\leq\operatorname{poly}(n/\varepsilon)$ and a primitive element for each field $\mathds{F}_{q}$.
The one-time preprocessing steps above can be implemented in time $\operatorname{polylog}(n/\varepsilon)$ using randomness. [^3] At least naively, deterministic algorithms for the above one-time preprocessing steps require time $\operatorname{poly}(n/\varepsilon)$, by testing primality of all numbers in an interval of length $\operatorname{poly}(n/\varepsilon)$. The preprocessing step in item 2 can be replaced by a different known procedure that can be implemented deterministically in time $\widetilde{O}(n+\sqrt{n}\log(1/\varepsilon)^{4+\delta})$ for any constant $\delta>0$. Note that this is $\widetilde{O}(n)$ when, say, $\varepsilon\geq 2^{-Cn^{0.1}}$, meaning that in this case its runtime can be absorbed into the runtime of the extractor, yielding item 1. This constraint can be weakened to roughly $\varepsilon\geq 2^{-Cn^{0.16}}$ by using an alternative approach that we sketch in section 3.3.3. To avoid overloading 1, we leave these discussions to remark 3.11 and section 3.3.3, respectively.
˜1 follows from combining modern condensers with short seeds (namely, the lossless condenser of Kalev and Ta-Shma [KT22] and the lossy Reed-Solomon-based condenser of Guruswami, Umans, and Vadhan [GUV09]) with a careful combination and instantiation of classical recursive approaches developed by Srinivasan and Zuckerman [SZ99] and in [GUV09]. It readily implies, among other things, an $\widetilde{O}(n)$-time privacy amplification protocol where only $O(\log(n/\varepsilon))$ bits need to be communicated over the one-way authenticated public channel and almost all the min-entropy can be extracted (after a reasonable one-time preprocessing step if $k$ or $\varepsilon$ are very small).
Remark 1.1 (complexity of the preprocessing steps).
Both of the one-time preprocessing steps in ˜1 are well-studied, and can be implemented in time $\operatorname{polylog}(n/\varepsilon)$ using randomness. They are related to the condensers we use in the construction, and we discuss their complexity in more detail in remarks˜3.11 and 3.16, and some approaches (and barriers) towards derandomizing them in section˜3.3.3.
@@
As a conceptual contribution which may be of independent interest, we present a new “non-recursive” construction of extractors with seed length $O(\log(n/\varepsilon))$ and output length $(1-\eta)k$ that is computable in nearly-linear time when $k>\operatorname{polylog}(1/\varepsilon)$ and avoids the complicated recursive procedures from [SZ99, GUV09] and ˜1. We believe this to be a conceptually better approach towards constructing seeded extractors, and we discuss it in more detail in the technical overview and fully in section˜4.1.
@@
One of the most widely-used explicit seeded extractors is Trevisan’s extractor [TRE01, RRV02]. While by now we have extractors with better parameters, one of its main advantages is that it is one of the few examples of extractors, and in a sense the best one, which are known to be quantum-proof. [^4] An extractor is quantum-proof if its output is close to uniform even in the presence of a quantum adversary that has some (bounded) correlation with $X$. A bit more formally, $\mathsf{Ext}$ is quantum-proof if for all classical-quantum states $\rho_{XE}$ (where $E$ is a quantum state correlated with $X$) with $H_{\infty}(X|E)\geq k$, and a uniform seed $Y$, it holds that $\rho_{\mathsf{Ext}(X,Y)YE}\approx_{\varepsilon}\rho_{U_{m}}\otimes\rho_{Y}\otimes\rho_{E}$. See [DPV+12] for more details.
Trevisan’s extractor uses two basic primitives: combinatorial designs (when more than one output bit is desired), and binary list-decodable codes. A standard instantiation of such suitable codes goes by concatenating a Reed-Solomon code with a Hadamard code, and this is also what is considered in [FWE+23, FYE+25]. As they also observe, this gives a nearly-linear time construction when the output length $m=1$. In fact, by leveraging fast multipoint evaluation, one can also get a nearly-linear time construction for any output length $m\leq\frac{n}{\log(1/\varepsilon)}$, although this was not noted in previous works. [^5] For a rigorous statement on fast multipoint evaluation, see lemma 2.1.
We present an alternative instantiation of Trevisan’s extractor that can be computed in truly linear time on a RAM in the logarithmic cost model, for any output length $m\leq\frac{n}{\log(1/\varepsilon)\cdot\operatorname{polylog}(n)}$. While the underlying technical details are simple, we opt to present this result here because it may be of interest to some readers due to wide interest in efficient implementations of Trevisan’s extractor, and because it was not observed in prior works.
@@
    Time $O(n)+m\log(1/\varepsilon)\cdot\operatorname{polylog}(n)$ after a preprocessing step [^6] This preprocessing step corresponds to precomputing the design, and is not the same preprocessing step as in 1. running in time $\widetilde{O}(m\log(n/\varepsilon))$, on a RAM in the logarithmic cost model. In particular, there exists a universal constant $c$, such that whenever $m\leq\frac{n}{\log(1/\varepsilon)\cdot\log^{c}(n)}$, the instantiation runs in time $O(n)$, without the need for a preprocessing step.
@@ 1.2 Other Related Work
Besides the long line of work focusing on improved constructions of explicit seeded extractors and mentioned in the introduction above, other works have studied randomness extraction in a variety of restricted computational models. These include extractors computable by streaming algorithms [BRS+02], local algorithms [LU02, VAD04, BG13, CL18], AC0 circuits [GVW15, CL18, CW24], AC0 circuits with a layer of parity gates [HIV22], NC1 circuits [CW24], and low-degree polynomials [ACG+22, AGM+25, GGH+24]. Moreover, some works have independently explored implementations of other fundamental pseudorandomness primitives in various restricted computational models. These include $k$-wise and $\varepsilon$-biased generators, which often play a key role in constructions of various types of extractors. See [HV06, 40, CRS+13, 50] for a very partial list.
As mentioned briefly above, some works have also focused on constructing seeded extractors computable in time $O(n\log n)$, motivated by applications in privacy amplification for quantum key distribution. Such constructions are based on hash functions, and are thus far restricted to $\Omega(m)$ seed length. The work of Hayashi and Tsurumaru [HT16] presents an extensive discussion of such efforts. We also mention that nearly-linear time extractors with very short seed, in the regime $k=n^{\Omega(1)}$ and $\varepsilon=n^{-o(1)}$, were given in [DMO+22], with applications in derandomization. [^7] Our extractor can replace the one constructed in [DMO+22], and it is indeed more efficient. However, due to other bottlenecks, they need to work with high error $\varepsilon=n^{-o(1)}$ and relatively large min-entropy, and so the difference between the two extractors is not significant.
The techniques in this paper build on, and extend, block-source extraction techniques [NZ96, ZUC96, ZUC97, SZ99, GUV09]. Another line of work, notably including [NT99, LRV+03, DW08, DKS+13, TU12], utilizes mergers to construct seeded extractors. [^8] For the definition of mergers, see, e.g., [DKS+13]. We note that those works do often use block source conversion techniques, but usually in a different manner. However, when restricted to constructions that get optimal seed length, they generally do not support low error (let alone run in nearly-linear time). On the positive side, the state-of-the-art mergers-based constructions get sub-linear entropy loss [DKS+13, TU12], whereas the constructions in this paper do not. We stress that it is still a very interesting open problem to construct a low-error, optimal seed length extractor with sub-linear entropy loss.